verify
(sinx+tanx)/(cosx+1)=tanx

Answer by Alan3354(30959)

(sinx+tanx)/(cosx+1)=tanx
sin + tan = cos*tan + tan
sin + tan = cos*(sin/cos) + tan
sin + tan = sin + tan